A 3.94 gram sample of oxygen gas has a volume of 937 milliliters at a pressure of 4.00 atm. The temperature of the O2 gas sample is  °C.
Answer #1

From ideal gas equation, PV = nRT

The number of moles of O2 gas = 3.94/32=0.123 moles

The temperature of the gas, T = PV/nR = 4*0.937/(0.123*0.0821) = 371.15 K

The temperature of the O2 gas in Celsius is

371.15-273 = 98.15 degree Celsius